Suddenly the new problem with our Agilent 7000A has appeared
Instrument has worked quite intensively for a year in our laboratory (instrument has been bought refurbished 20 month ago).
A month ago, during the analysis process, error 1.5 "Entrance lens driver cannot maintain the requested voltage" began to appear. After some time, the error is cleared and several tens of seconds it is OK, and then the same error 1.5 appears again.
At the same time the high vacuum sensor begins to give overestimated readings: normally the vacuum is ~ 5x10e-5 Torr, and at the moment when an error appears, the sensor gives 10e-8 - 10e-10 Torr on the MS display, and Mass Hunter shows only ---. When everything comes back to normal operation, the readings on the MS display and in the MH become equal and have the order of 5x10e-5
When we tried to make a ramp in Manual Tune (Ion source/Entrance lens), there was no power to the filaments. This coincided in time with incorrect readings of the high vacuum sensor. When the high vacuum sensor readings became normal, the ramp was acquired normally
What has been done:
1). Ion source has been cleaned
2). Filaments were replaced (the old ones looked good).
3). Checked the contacts of the ion source board
4). Checked lens-body short circuit - OK
5). Wires and pins - OK
6). There is no shorts between pins and body
7). Sideboards were cleaned with compressed air and vacuum cleaner (there was quite a lot of dust)
According to the service engineer, the problem was the high vacuum sensor. The high vacuum sensor gives incorrect readings that are outside the acceptable range, and the protection system turns off the power of the ion source - hence the error 1.5 (electrical supply - Entrance lens.
The high vacuum sensor has been ordered and replaced. For about five minutes after starting the MS (turbine at 100%) the high vacuum values were normal. At this moment, we acquired correct IS/Entrance lens ramp in Manual Tune and after that error 1.5 "Entrance lens driver cannot maintain the requested voltage" appeared again, the readings of the high vacuum sensor went to the 10e-8 - 10e-10 Torr unrealistic range.
Analyzing LogBook we also revealed that along with error 1.5 an error 5.1 “Firmware DSP error” also appeared periodically
What could be source of the malfunction? What else can be diagnosed?